Find two consecutive negative odd integers whose product is 399
Answer by TimothyLamb(4379)

---
x - y = -2
x = y - 2
xy = 399
---
xy = 399
(y - 2)y = 399
yy - 2y - 399 = 0
---
the above quadratic equation is in standard form, with a=1, b=-2, and c=-399
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -2 -399
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
y = 21
y = -19
---
positive root doesn't satisfy the problem statement, so use the negative root:
---
answer:
y = -19
x = -21
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php